### v1

@@ Abstract
Foundation models pretrained on large-scale datasets are revolutionizing the field of computational pathology (CPath). The generalization ability of foundation models is crucial for the success in various downstream clinical tasks. However, current foundation models have only been evaluated on a limited type and number of tasks, leaving their generalization ability and overall performance unclear. To address this gap, we established a most comprehensive benchmark to evaluate the performance of off-the-shelf foundation models across six distinct clinical task types, encompassing a total of 39 specific tasks. Our findings reveal that existing foundation models excel at certain task types but struggle to effectively handle the full breadth of clinical tasks. To improve the generalization of pathology foundation models, we propose a unified knowledge distillation framework consisting of both expert and self knowledge distillation, where the former allows the model to learn from the knowledge of multiple expert models, while the latter leverages self-distillation to enable image representation learning via local-global alignment. Based on this framework, a Generalizable Pathology Foundation Model (GPFM) is pretrained on a large-scale dataset consisting of 190 million images from around 86,000 public H&E whole slides across 34 major tissue types. Evaluated on the established benchmark, GPFM achieves an impressive average rank of 1.36, with 29 tasks ranked 1st, while the the second-best model, UNI, attains an average rank of 2.96, with only 4 tasks ranked 1st. The superior generalization of GPFM demonstrates its exceptional modeling capabilities across a wide range of clinical tasks, positioning it as a new cornerstone for feature representation in CPath.
@@ 1 Introduction
In recent years, there has been a notable progress in the fields of computer vision and natural language processing driven by self-supervised learning on large-scale datasets. These pretrained models, commonly referred to as foundation models, have garnered significant attention and have exhibited remarkable success across various tasks [28, 29, 30]. In the field of CPath, some efforts [31, 32, 33, 34, 35, 36, 37] have been dedicated to pretraining foundation models that can learn inherent representations of histopathology images, catering to the diverse array of tasks encountered in clinical pathology practice. However, the current foundation models have only been evaluated on a limited type of tasks as shown in Figure 2.a, leaving their overall performance unclear. To comprehensively evaluate these models, we built a most comprehensive benchmark spanning six major clinical task categories, comprising 39 specific tasks. Our findings revealed that the generalization ability of these models is still limited and no single model can effectively address all the tasks as shown in Figure 1.c. It can be seen that UNI [33] achieves best performance in image retrieval and patch-level (ROI) tissues classification tasks, Phikon [32] performs best in survival analysis and report generation tasks, and CONCH [35] obtains highest performance in WSI classification and visual question answering (VQA) tasks. This can be attributed to the fact that each foundation model is trained using distinct datasets and pretraining strategies, leading to specific advantages for each model within particular datasets. These findings highlight the need for further research to develop more generalizable foundation models that can consistently perform well across the diverse type of clinical tasks encountered in CPath. By addressing this challenge, we can unlock the full potential of foundation model in CPath.
@@
To improve the generalization of pathology foundation model and enhance the overall performance, an intuitive idea is to leverage the specific strengths of existing models by employing knowledge distillation techniques [38, 39]. Accordingly, we proposed a novel self-supervised learning framework with expert and self knowledge distillation to develop a Generalizable Pathology Foundation Model (GPFM). Based on this pretraining method, we trained GPFM using a diverse dataset of approximately 190 million images from 86,104 public WSIs across 34 major tissue types. The overview of GPFM is presented in Figure 1. With the collected diverse tissues and the indirectly using of the images that used to pretrain expert models (e.g., UNI and CONCH), GPFM exhibits outstanding performance across the established benchmarks, achieving average rank of 1.36 (29 out of 39 tasks), while the second-best performing model, UNI, achieving average rank of 2.96 (4 out of 39 tasks). These results demonstrate the efficacy of GPFM as a generalizable foundation model in CPath, showcasing its potential to significantly advance the field. The ability of GPFM to consistently perform well across a diverse type of clinical tasks highlights the benefits of leveraging knowledge distillation to combine the strengths of expert models, ultimately leading to more robust and versatile foundation models for supporting clinicians and improving patient care.
@@ 2.1 WSI Classification
Figure 3: Performance of Foundation Models on WSI Classification Tasks. a-c. Average AUC, balanced accuracy (ACC), and weighted F1 score of foundation models across 12 WSI classification tasks. d. Average ranking order of foundation models based on AUC across 12 WSI classification tasks. e-l. The AUC scores of foundation models on 8 WSI classification tasks. The violin plots depict the distribution of results based on 1,000 bootstrap replicates. The Wilcoxon signed-rank one-sided test was used to detect significant differences between the top-performing model (GPFM) and the model with the next-best performance, when applicable. Then horizontal black line in the violin plot is the mean AUC. The error bar indicates the 95% confidence intervals. It is worth noting that the WSIs of BRCA and RCC are from TCGA. More results are represented in Extended Data Figure 9.
@@
Across 12 WSI classification tasks, GPFM achieved an average ranking score of 1.08, outperforming the second-best performing model, UNI, which achieves an average ranking score of 3.04. It’s worth noting that the ranking is based on the Area Under the Curve (AUC) metric, where a higher AUC indicates better performance. The detailed ranking scores of WSI classification tasks are presented in Figure 2.d. Furthermore, we evaluated the overall performance using average metrics, including AUC, balanced accuracy, and weighted F1 score. As shown in Figure 3.a, GPFM achieved the best average AUC of 0.956, representing a 1.4% improvement over the second-best performing model, CONCH, with statistical significance P < 0.001 (+1.4%, CONCH, P<0.001). Similarly, as depicted in Figure 3.b-c, GPFM achieved the best balanced accuracy of 0.833 (+1.9%, UNI, P<0.001) and the best weighted F1 score of 0.834 (+1.9%, UNI, P<0.001). The detailed performance of the different foundation models are presented in Figure 3.e-l and Extended Data Figure 9.a-d. We also visualized the ROC curves of GPFM in Extended Data Figure10 for every classification task. These results across multiple performance metrics clearly illustrated the great potential and strong generalization ability of the GPFM model in the WSI classification tasks.
@@
GPFM advances gene mutation prediction. Follow same evaluation strategy as cancer diagnosis, we also investigated the potential of foundation models in 2 gene mutation prediction tasks. The GPFM achieved the best performance both in LUAD-TP53 (from TCGA) and Glioma IDH 1 (from TCGA GBMLGG) tasks, with AUCs of 0.855 and 0.998, respectively. Compared to the second-best model, Phikon, the GPFM showed a 1.2% improvement in AUC. These results illustrate that the GPFM is more generalizable than existing foundation models for gene mutation prediction tasks.
@@ 2.3 ROI Classification
Across all 12 ROI classification tasks, GPFM achieved the best ranking score of 1.60, while the second-best performing model, UNI, attained ranking score of 2.67. The detailed ranking socres are presented in Figure 2.d. Furthermore, when considering the conventional metrics, GPFM attained the highest AUC of 0.955 (+0.2%, UNI, P<0.001), the best weighted F1 score of 0.848 (+0.8%, UNI, P<0.001), and the top balanced accuracy of 0.853 (+0.8%, UNI, P<0.001) on average. The comparison of various foundation models is illustrated in Figure 5. Specifically, GPFM demonstrated the best performance in the CRC-MSI, PanCancer-TIL, PCAM, WSSS4LUAD datasets, with AUC values of 0.812 (+0.2%, CONCH, P<0.001), 0.978 (+0.1%, UNI, P<0.001), 0.988 (+0.6%, UNI, P<0.001), 0.998 (+0.1%, Phikon, P<0.001), respectively. Since ROI classification tasks are relatively easy, GPFM also jointly ranked first with other foundation models in the CRC-100K, CCRCC-TCGA-HEL, and PanCancer-TCGA datasets, achieving AUC values of 0.995 (Ctranspath), 0.997 ( Phikon), 0.999 (Phikon), respectively. We also conducted external validation on Chaoyang, UniToPatho, and BreakHis datasets. Across the three external validation datasets, the GPFM achieved the best average rank of 1.5, while the second best performing model, UNI, obtained an average rank of 2.3, illustrating the generalization capability of GPFM.
@@
An examination of the average ranking score in Figure 5.a reveals that the performance of Phikon, CONCH and UNI perform closely, with ranking score differences less than 1. This suggests that no single existing model dominates in ROI classification tasks. However, GPFM achieves an averaging ranking score of 1.60, more than one point ahead of the next best model, demonstrating its strength as a generalizable foundation model. The above results also verify the effectiveness of the proposed unified knowledge distillation approach. By leveraging the knowledge of all adopted expert models, the unified knowledge distillation method is able to outperform the individual expert model.
@@ 2.4 Pathological Image Retrieval
Image retrieval technique could match the new patient pathology images to a curated database of previously diagnosed cases, providing pathologists with a novel tool to enhance diagnostic accuracy. Through visual inspection and comparison of similar historical cases, pathologists can leverage image search functionality to enhance their diagnostic decision-making. This approach allows pathologists to go beyond the capabilities of individual AI algorithms and instead harness the collective knowledge distilled from large repositories of validated pathology images and expert diagnoses [57, 58, 59]. In this study, we employ the CRC-100K dataset [60] for conducting pathological image retrieval tasks. This dataset comprises 100,000 non-overlapping 224$\times$224 patches extracted from 86 human cancer tissue slides stained with H&E for training purposes. Additionally, it includes 7,180 images with 224$\times$224 pixels extracted from 50 patients diagnosed with colorectal adenocarcinoma for testing. The dataset consists of multiple classes, including Adipose (ADI, 11,745 ROIs), Background (BACK, 11,413 ROIs), Debris (DEB, 11,851 ROIs), Lymphocytes (LYM, 12,191 ROIs), Mucus (MUC, 9,931 ROIs), Smooth muscle (MUS, 14,128 ROIs), Normal colon mucosa (NORM, 9,504 ROIs), Cancer-associated stroma (STR, 10,867 ROIs), and Colorectal adenocarcinoma epithelium (TUM, 15,550 ROIs). For training and evaluation, we utilize the official train-test split, with 100,000 samples for training and 7,180 samples for testing.
@@ 2.6 Pathology Report Generation
Pathology reports are essential components of the healthcare system, providing critical information to clinicians and patients about the diagnosis, prognosis, and treatment of various medical conditions. These reports summarize the findings from pathological examinations, such as biopsies, cytology samples, and surgical specimens, and play a vital role in guiding clinical decision-making. Traditionally, pathology reports are written manually by pathologists and their teams, a time-consuming and labor-intensive process. Recent advancements in natural language processing (NLP) and machine learning have enabled the development of automated pathology report generation systems, which can dramatically improve the efficiency and consistency of this critical task [63, 64, 65]. To evaluate the effectiveness of foundation models in pathology report generation, we utilize the TCGA WSI-report dataset curated by Guo et al. [63]. This dataset consists of 7,690 pairs of WSIs and their corresponding diagnosis reports, covering various diseases from different primary sites. In the task of WSI report generation, foundation models are employed to provide representations of WSIs as inputs to the report generation model. We use HistGen [63] as the baseline model, which takes the WSI features extracted by foundation models as input and generates the corresponding diagnosis report. For training and evaluation, we split the dataset into train-validation-test cohorts with a ratio of 7:1:2 (5,383: 769: 1,538 slides). The performance of each foundation model in report generation is presented in Figure 7 and the detailed results is reported in Extended Data Table 35.
@@
The experimental results demonstrate that Phikon achieved the best performance across all six metrics, while GPFM achieved comparable performance and ranked as the second-best model. This is quite surprising that vision foundation models (e.g., Phikon and GPFM) achieved much better performance in this task compared to vision-language foundation models such as CONCH and PLIP. A potential reason behind is that PLIP and CONCH were only trained with short descriptions or captions of pathological images, without access to the global contextual information. This may have made the text-image pairs less effective for this task compared to the VQA tasks they were originally designed for. It is worth noting that the Phikon model was pretrained exclusively on data from TCGA. While this may enhance Phikon’s efficacy on tasks specifically utilizing data derived from the TCGA, it also limited Phikon’s generalization ability on tasks outside of the TCGA domain. To leverage the complementary strengths of existing models, the proposed unified knowledge distillation approach can distill the capabilities of Phikon in report generation into the GPFM. This synergistic integration allows us to combine the respective strengths of these foundation models, leading to the development of a more generalizable model.
@@
Figure 7: Performance of Report Generation. a. The performance of report generation measured by six different metrics. b. Examples of generated pathology reports. The ground truth reports provided by pathologists are shown for comparison. The text in red indicates correct predictions, the text in blue represents incorrect predictions. The reports are formatted without changing the original content. More examples are represented in Extended Data Figure 13 and Figure 14.
@@ 2.7 The Effectiveness of Expert Knowledge Distillation
In the self-supervised learning framework proposed in this study, we introduced a unified knowledge distillation model to facilitate the transfer of knowledge from off-the-shelf foundation models to GPFM during the pretraining stage. To assess the effectiveness of this module, we conducted an experiment where we removed the Expert Knowledge Distillation module, resulting in a modified self-supervised learning framework known as DINOv2 [56]. We trained both DINOv2 and GPFM on the same dataset and evaluated their performance in tissue classification tasks. The experimental results are presented in Figure 8. More details can be found in Extended Data Table 36. The experimental results clearly demonstrate the positive impact of Expert Knowledge Distillation on the performance of the models across 12 tasks. Not only did the individual task performances improve significantly, but the average performance also exhibited enhancement, with notable improvements in all three metrics. The AUC increased by 0.6%, the weighted F1 score improved by 1.8%, and the balanced accuracy showed an increase of 1.8%. These findings provide strong evidence for the effectiveness of transferring the knowledge from off-the-shelf pathology foundation models through the proposed knowledge distillation learning framework.
Figure 8: The Effectiveness of Expert Knowledge Distillation. The figure presents the performance difference between GPFM (with Expert Knowledge Distillation) and DINOv2 (without Expert Knowledge Distillation). The horizontal black lines indicate mean AUC. If GPFM outperforms DINOv2, the p-value is also reported. a. The balanced accuracy of the models with and without expert knowledge distillation. b. The weighted F1 score of the models with and without expert knowledge distillation. c. The AUC of the models with and without expert knowledge distillation. Significance testing was conducted using the Wilcoxon signed-rank one-sided test, demonstrating that expert knowledge distillation consistently improves performance across the majority of tasks, highlighting the effectiveness of this technique in enhancing the GPFM.

### v2

@@ Abstract
Foundation models pretrained on large-scale datasets are revolutionizing the field of computational pathology (CPath). The generalization ability of foundation models is crucial for the success in various downstream clinical tasks. However, current foundation models have only been evaluated on a limited type and number of tasks, leaving their generalization ability and overall performance unclear. To address this gap, we established a most comprehensive benchmark to evaluate the performance of off-the-shelf foundation models across six distinct clinical task types, encompassing a total of 39 specific tasks. Our findings reveal that existing foundation models excel at certain task types but struggle to effectively handle the full breadth of clinical tasks. To improve the generalization of pathology foundation models, we propose a unified knowledge distillation framework consisting of both expert and self-knowledge distillation, where the former allows the model to learn from the knowledge of multiple expert models, while the latter leverages self-distillation to enable image representation learning via local-global alignment. Based on this framework, a Generalizable Pathology Foundation Model (GPFM) is pretrained on a large-scale dataset consisting of 190 million images from around 86,000 public H&E whole slides across 34 major tissue types. Evaluated on the established benchmark, GPFM achieves an impressive average rank of 1.36, with 29 tasks ranked 1st, while the second-best model, UNI, attains an average rank of 2.96, with only 4 tasks ranked 1st. The superior generalization of GPFM demonstrates its exceptional modeling capabilities across a wide range of clinical tasks, positioning it as a new cornerstone for feature representation in CPath.
@@ 1 Introduction
In recent years, there has been a notable progress in the fields of computer vision and natural language processing driven by self-supervised learning on large-scale datasets. These pretrained models, commonly referred to as foundation models, have garnered significant attention and have exhibited remarkable success across various tasks [28, 29, 30]. In the field of CPath, some efforts [31, 32, 33, 34, 35, 36, 37] have been dedicated to pretraining foundation models that can learn inherent representations of histopathology images, catering to the diverse array of tasks encountered in clinical pathology practice. However, the current foundation models have only been evaluated on a limited type of tasks as shown in Figure 2.a, leaving their overall performance unclear. To comprehensively evaluate these models, we built a most comprehensive benchmark spanning six major clinical task categories, comprising 39 specific tasks. Our findings revealed that the generalization ability of these models is still limited and no single model can effectively address all the tasks as shown in Figure 1.c. It can be seen that UNI [33] achieves the best performance in image retrieval and patch-level (ROI) tissue classification tasks, Phikon [32] performs best in survival analysis and report generation tasks, and CONCH [35] obtains highest performance in WSI classification and visual question answering (VQA) tasks. This can be attributed to the fact that each foundation model is trained using distinct datasets and pretraining strategies, leading to specific advantages for each model within particular datasets. These findings highlight the need for further research to develop more generalizable foundation models that can consistently perform well across the diverse type of clinical tasks encountered in CPath. By addressing this challenge, we can unlock the full potential of the foundation model in CPath.
@@
To improve the generalization of pathology foundation model and enhance the overall performance, an intuitive idea is to leverage the specific strengths of existing models by employing knowledge distillation techniques [38, 39]. Accordingly, we proposed a novel self-supervised learning framework with expert and self knowledge distillation to develop a Generalizable Pathology Foundation Model (GPFM). Based on this pretraining method, we trained GPFM using a diverse dataset of approximately 190 million images from 86,104 public WSIs across 34 major tissue types. The overview of GPFM is presented in Figure 1. With the collected diverse tissues and the indirectly using of the images that used to pretrain expert models (e.g., UNI and CONCH), GPFM exhibits outstanding performance across the established benchmarks, achieving an average rank of 1.36 (29 out of 39 tasks), while the second-best performing model, UNI, achieving an average rank of 2.96 (4 out of 39 tasks). These results demonstrate the efficacy of GPFM as a generalizable foundation model in CPath, showcasing its potential to significantly advance the field. The ability of GPFM to consistently perform well across a diverse type of clinical tasks highlights the benefits of leveraging knowledge distillation to combine the strengths of expert models, ultimately leading to more robust and versatile foundation models for supporting clinicians and improving patient care.
@@ 2.1 WSI Classification
Figure 3: Performance of Foundation Models on WSI Classification Tasks. a-c. Average AUC, balanced accuracy (ACC), and weighted F1 score of foundation models across 12 WSI classification tasks. d. Average ranking order of foundation models based on AUC across 12 WSI classification tasks. e-l. The AUC scores of foundation models on 8 WSI classification tasks. The violin plots depict the distribution of results based on 1,000 bootstrap replicates. The Wilcoxon signed-rank one-sided test was used to detect significant differences between the top-performing model (GPFM) and the model with the next-best performance, when applicable. Then horizontal black line in the violin plot is the mean AUC. The error bar indicates the 95% confidence interval. It is worth noting that the WSIs of BRCA and RCC are from TCGA. More results are represented in Extended Data Figure 9.
@@
Across 12 WSI classification tasks, GPFM achieved an average ranking score of 1.08, outperforming the second-best performing model, UNI, which achieved an average ranking score of 3.04. It’s worth noting that the ranking is based on the Area Under the Curve (AUC) metric, where a higher AUC indicates better performance. The detailed ranking scores of WSI classification tasks are presented in Figure 2.d. Furthermore, we evaluated the overall performance using average metrics, including AUC, balanced accuracy, and weighted F1 score. As shown in Figure 3.a, GPFM achieved the best average AUC of 0.956, representing a 1.4% improvement over the second-best performing model, CONCH, with statistical significance P < 0.001 (+1.4%, CONCH, P<0.001). Similarly, as depicted in Figure 3.b-c, GPFM achieved the best balanced accuracy of 0.833 (+1.9%, UNI, P<0.001) and the best weighted F1 score of 0.834 (+1.9%, UNI, P<0.001). The detailed performance of the different foundation models is presented in Figure 3.e-l and Extended Data Figure 9.a-d. We also visualized the ROC curves of GPFM in Extended Data Figure10 for every classification task. These results across multiple performance metrics clearly illustrated the great potential and strong generalization ability of the GPFM model in the WSI classification tasks.
@@
GPFM advances gene mutation prediction. Following the same evaluation strategy as cancer diagnosis, we also investigated the potential of foundation models in 2 gene mutation prediction tasks. The GPFM achieved the best performance both in LUAD-TP53 (from TCGA) and Glioma IDH 1 (from TCGA GBMLGG) tasks, with AUCs of 0.855 and 0.998, respectively. Compared to the second-best model, Phikon, the GPFM showed a 1.2% improvement in AUC. These results illustrate that the GPFM is more generalizable than existing foundation models for gene mutation prediction tasks.
@@ 2.3 ROI Classification
Across all 12 ROI classification tasks, GPFM achieved the best ranking score of 1.60, while the second-best performing model, UNI, attained a ranking score of 2.67. The detailed ranking scores are presented in Figure 2.d. Furthermore, when considering the conventional metrics, GPFM attained the highest AUC of 0.955 (+0.2%, UNI, P<0.001), the best weighted F1 score of 0.848 (+0.8%, UNI, P<0.001), and the top balanced accuracy of 0.853 (+0.8%, UNI, P<0.001) on average. The comparison of various foundation models is illustrated in Figure 5. Specifically, GPFM demonstrated the best performance in the CRC-MSI, PanCancer-TIL, PCAM, WSSS4LUAD datasets, with AUC values of 0.812 (+0.2%, CONCH, P<0.001), 0.978 (+0.1%, UNI, P<0.001), 0.988 (+0.6%, UNI, P<0.001), 0.998 (+0.1%, Phikon, P<0.001), respectively. Since ROI classification tasks are relatively easy, GPFM also jointly ranked first with other foundation models in the CRC-100K, CCRCC-TCGA-HEL, and PanCancer-TCGA datasets, achieving AUC values of 0.995 (Ctranspath), 0.997 ( Phikon), 0.999 (Phikon), respectively. We also conducted external validation on Chaoyang, UniToPatho, and BreakHis datasets. Across the three external validation datasets, the GPFM achieved the best average rank of 1.5, while the second best performing model, UNI, obtained an average rank of 2.3, illustrating the generalization capability of GPFM.
@@
An examination of the average ranking score in Figure 5.a reveals that the performance of Phikon, CONCH, and UNI perform closely, with ranking score differences less than 1. This suggests that no single existing model dominates in ROI classification tasks. However, GPFM achieves an average ranking score of 1.60, more than one point ahead of the next best model, demonstrating its strength as a generalizable foundation model. The above results also verify the effectiveness of the proposed unified knowledge distillation approach. By leveraging the knowledge of all adopted expert models, the unified knowledge distillation method is able to outperform the individual expert model.
@@ 2.4 Pathological Image Retrieval
Image retrieval techniques could match the new patient pathology images to a curated database of previously diagnosed cases, providing pathologists with a novel tool to enhance diagnostic accuracy. Through visual inspection and comparison of similar historical cases, pathologists can leverage image search functionality to enhance their diagnostic decision-making. This approach allows pathologists to go beyond the capabilities of individual AI algorithms and instead harness the collective knowledge distilled from large repositories of validated pathology images and expert diagnoses [57, 58, 59]. In this study, we employ the CRC-100K dataset [60] for conducting pathological image retrieval tasks. This dataset comprises 100,000 non-overlapping 224$\times$224 patches extracted from 86 human cancer tissue slides stained with H&E for training purposes. Additionally, it includes 7,180 images with 224$\times$224 pixels extracted from 50 patients diagnosed with colorectal adenocarcinoma for testing. The dataset consists of multiple classes, including Adipose (ADI, 11,745 ROIs), Background (BACK, 11,413 ROIs), Debris (DEB, 11,851 ROIs), Lymphocytes (LYM, 12,191 ROIs), Mucus (MUC, 9,931 ROIs), Smooth muscle (MUS, 14,128 ROIs), Normal colon mucosa (NORM, 9,504 ROIs), Cancer-associated stroma (STR, 10,867 ROIs), and Colorectal adenocarcinoma epithelium (TUM, 15,550 ROIs). For training and evaluation, we utilize the official train-test split, with 100,000 samples for training and 7,180 samples for testing.
@@ 2.6 Pathology Report Generation
Pathology reports are essential components of the healthcare system, providing critical information to clinicians and patients about the diagnosis, prognosis, and treatment of various medical conditions. These reports summarize the findings from pathological examinations, such as biopsies, cytology samples, and surgical specimens, and play a vital role in guiding clinical decision-making. Traditionally, pathology reports are written manually by pathologists and their teams, a time-consuming and labor-intensive process. Recent advancements in natural language processing (NLP) and machine learning have enabled the development of automated pathology report generation systems, which can dramatically improve the efficiency and consistency of this critical task [63, 64, 65]. To evaluate the effectiveness of foundation models in pathology report generation, we utilize the TCGA WSI-report dataset curated by Guo et al. [63]. This dataset consists of 7,690 pairs of WSIs and their corresponding diagnosis reports, covering various diseases from different primary sites. In the task of WSI report generation, foundation models are employed to provide representations of WSIs as inputs to the report generation model. We use HistGen [63] as the baseline model, which takes the WSI features extracted by foundation models as input and generates the corresponding diagnosis report. For training and evaluation, we split the dataset into train-validation-test cohorts with a ratio of 7:1:2 (5,383: 769: 1,538 slides). The performance of each foundation model in report generation is presented in Figure 7 and the detailed results are reported in Extended Data Table 35.
@@
The experimental results demonstrate that Phikon achieved the best performance across all six metrics, while GPFM achieved comparable performance and ranked as the second-best model. This is quite surprising that vision foundation models (e.g., Phikon and GPFM) achieved much better performance in this task compared to vision-language foundation models such as CONCH and PLIP. A potential reason behind is that PLIP and CONCH were only trained with short descriptions or captions of pathological images, without access to the global contextual information. This may have made the text-image pairs less effective for this task compared to the VQA tasks they were originally designed for. It is worth noting that the Phikon model was pretrained exclusively on data from TCGA. While this may enhance Phikon’s efficacy on tasks specifically utilizing data derived from the TCGA, it also limits Phikon’s generalization ability on tasks outside of the TCGA domain. To leverage the complementary strengths of existing models, the proposed unified knowledge distillation approach can distill the capabilities of Phikon in report generation into the GPFM. This synergistic integration allows us to combine the respective strengths of these foundation models, leading to the development of a more generalizable model.
@@
Figure 7: Performance of Report Generation. a. The performance of report generation is measured by six different metrics. b. Examples of generated pathology reports. The ground truth reports provided by pathologists are shown for comparison. The text in red indicates correct predictions, and the text in blue represents incorrect predictions. The reports are formatted without changing the original content. More examples are represented in Extended Data Figure 13 and Figure 14.
@@ 2.7 The Effectiveness of Expert Knowledge Distillation
In the self-supervised learning framework proposed in this study, we introduced a unified knowledge distillation model to facilitate the transfer of knowledge from off-the-shelf foundation models to GPFM during the pretraining stage. To assess the effectiveness of this module, we conducted an experiment where we removed the Expert Knowledge Distillation module, resulting in a modified self-supervised learning framework known as DINOv2 [56]. We trained both DINOv2 and GPFM on the same dataset and evaluated their performance in tissue classification tasks. The experimental results are presented in Figure 8. More details can be found in Extended Data Table 36. The experimental results clearly demonstrate the positive impact of Expert Knowledge Distillation on the performance of the models across 12 tasks. Not only did the individual task performances improve significantly, but the average performance also exhibited enhancement, with notable improvements in all three metrics. The AUC increased by 0.6%, the weighted F1 score improved by 1.8%, and the balanced accuracy showed an increase of 1.8%. These findings provide strong evidence for the effectiveness of transferring knowledge from off-the-shelf pathology foundation models through the proposed knowledge distillation learning framework.
Figure 8: The Effectiveness of Expert Knowledge Distillation. The figure presents the performance difference between GPFM (with Expert Knowledge Distillation) and DINOv2 (without Expert Knowledge Distillation). The horizontal black lines indicate the mean AUC. If GPFM outperforms DINOv2, the p-value is also reported. a. The balanced accuracy of the models with and without expert knowledge distillation. b. The weighted F1 score of the models with and without expert knowledge distillation. c. The AUC of the models with and without expert knowledge distillation. Significance testing was conducted using the Wilcoxon signed-rank one-sided test, demonstrating that expert knowledge distillation consistently improves performance across the majority of tasks, highlighting the effectiveness of this technique in enhancing the GPFM.